### v2

# envname-P envname#1

Abstract.
In this paper, we introduce the novel task of Open-domain Urban Itinerary Planning (OUIP), a paradigm designed to generate personalized urban itineraries from user requests articulated in natural language. This approach is different from traditional itinerary planning, which often restricts the granularity of user inputs, thus hindering genuine personalization. To this end, we present ItiNera, an OUIP system that synergizes spatial optimization with large language models (LLMs) to provide services that customize urban itineraries based on users’ needs. Upon receiving the user’s itinerary request, the LLM first decomposes it into detailed components, identifying key requirements, including preferences and dislikes. Then, we use these specifics to select candidate POIs from a large-scale collection using embedding-based Preference-aware POI Retrieval. Finally, a preference score-based Cluster-aware Spatial Optimization module clusters, filters, and orders these POIs, followed by the LLM for detailed POI selection and organization to craft a personalized, spatially coherent itinerary. Moreover, we created an LLM-based pipeline to update and personalize a user-owned POI database. This ensures up-to-date POI information, supports itinerary planning, pre-trip research, POI collection, recommendations, and more. To the best of our knowledge, this study marks the first integration of LLMs to innovate itinerary planning, with potential extensions for various urban travel and exploration activities. Offline and online evaluations demonstrate the capacity of our system to deliver more responsive, personalized, and spatially coherent itineraries than current solutions. Our system, deployed on an online platform, has attracted thousands of users for their urban travel planning.
Itinerary Planning; Intelligent Transportation; Location-based Services; Spatial Optimizations; Large Language Models
1. Introduction
As a novel form of urban travel and exploration, city tours invite travelers to wander through city streets, delve into historical sites, and immerse themselves in local culture. Hence it offers a more efficient, dynamic, and cost-effective travel experience compared to traditional tourism. Importantly, COVID-19 shifted travel trends towards crowd-free, health-conscious activities deeply connected with local contexts, propelling city tours to global popularity (Germano, 2023). However, planning a city tour is essentially an urban itinerary planning problem (Halder et al., 2024), a complex procedure that involves gathering travel-related information, selecting POIs, mapping out routes, and customizing to accommodate diverse user needs. An enjoyable city tour typically requires a significant time commitment in preparation.
Existing itinerary planning studies focus on traditional tourism. They typically consider coarse-grained user requirements, such as must-see POIs (Taylor et al., 2018), categories(Bolzoni et al., 2014), geographical considerations (Rani et al., 2018, 2019), and time budget (Zhang and Tang, 2018; Hsueh and Huang, 2019), or a combination of them (Yochum et al., 2020) as constraints to improve the logical arrangement of sites and optimize the quality of an itinerary (Chen et al., 2013; Sylejmani et al., 2017). We refer to this approach as spatial optimization, and it can maintain the quality of POIs and spatial coherency of the planned itinerary. However, the spatial optimization approach exhibits a lack of personalization and fails to ensure overall quality, thereby failing to satisfy the demands of users. A high-quality itinerary for a city tour should transcend a mere list of POIs, and importantly, it calls for alignment with users’ specific requirements and lifestyles.
Figure 1. The OUIP problem and the deployed OUIP service.
To elaborate on this, we summarize the highlighting features of city tours that differentiate itself from traditional tourism:
  • Dynamic Information: Unlike conventional tourism, where POIs remain relatively static, city tours involve rapidly changing POIs. This requires up-to-date information and allows for the exploration of temporary events, pop-up venues, and other fleeting experiences that traditional itineraries might overlook.
  • Individually Tailored: The planning of a city tour places significant emphasis on personal tastes and interests. Unlike traditional tourism, which often focuses on universally recognized POIs, city tours prioritize the unique preferences of the individual, making each itinerary deeply personalized.
  • Fine-grained Exploration: City tours encourage a detailed exploration of urban environments and a deeper understanding and appreciation of the city’s culture, architecture, and history, which goes beyond the surface-level glimpse offered by traditional tourism.
  • Diverse Constraints: City tour planning takes into account more complex constraints, necessitating more intricate planning. This includes implicit constraints that align with a person’s lifestyle and explicit constraints such as specific traveler interests or accessibility requirements.
For example, an ideal itinerary might start at a coffee shop in the morning, traverse through selected preferred POIs without detours, and conclude at an affordable yet atmospheric restaurant. An itinerary of this granularity not only fulfills travelers’ distinct preferences but significantly enriches the travel experience.
One can see that it is challenging to directly apply conventional itinerary planning approaches due to the above-discussed features of city tours. Ideally, users would articulate their preferences in a single sentence and, in return, receive a personalized, spatially coherent, well-designed itinerary that satisfies their requirements. Based on these observations, there is a need for a new itinerary planning paradigm that can integrate detailed user preferences to craft personalized and fine-grained travel experiences. In view of these, we propose the new task of open-domain urban itinerary planning (OUIP) in Definition 1.
Definition 1 (OUIP-layman’s version).
Open-domain Urban Itinerary Planning (OUIP) aims to generate personalized urban travel itineraries based on user preferences in natural language.
The term “open-domain” involves generating responses to queries across a wide range of topics without domain-specific constraints. Similarly, our proposed OUIP problem aims to generate personalized itineraries by interpreting users’ preferences expressed in natural language, as depicted in Fig. 1.
Recently, LLMs have demonstrated impressive abilities in language comprehension and instruction-following, enabling them to understand user needs, handle simple planning tasks, and generate responses based on their knowledge (OpenAI, [n. d.]). Recent trends propose employing LLMs for itinerary planning problems (Xie et al., 2024). However, the limitations of LLMs become apparent in itinerary planning due to reliance on non-real-time information, incomplete knowledge, and a lack of spatial awareness. Consequently, they are unable to independently deliver a satisfactory user experience in itinerary planning. The itineraries generated directly by LLMs could be circuitous, lack of detail, and include hallucinated, non-existent POIs, rendering them less practical for actual use.
Beyond the methodological gaps, evaluating various approaches to OUIP presents its own set of difficulties. First of all, there is a general lack of datasets for evaluating the OUIP approaches, and the true city tour plans are usually presented in the format of unstructured texts and images. The baseline methods for OUIP are also limited. More critically, considering the open-domain nature of the OUIP task, it is insufficient to rely solely on traditional rule-based metrics for evaluation.
Given above challenges, to address the task of OUIP, this paper proposes ItiNera, a holistic OUIP system that synergizes spatial optimization with LLMs. Specifically, ItiNera features a User-owned POI Collection module capable of automatically scraping and extracting POI features from web content, dynamically updating a user-owned, personalized POI database to ensure both the customization and timeliness of POI information, laying the groundwork for subsequent OUIP tasks. ItiNera consists of five modules: User-owned POI Collection, Request Decomposition, Preference-aware POI Retrieval, Cluster-aware Spatial Optimization, and Itinerary Generation. All five modules are LLM-assisted, and details will be presented later. Importantly, as the OUIP is a novel task, we collect and build the validation dataset and develop a series of metrics for a comprehensive evaluation framework. Lastly, ItiNera has been deployed as a real-world service that attracts thousands of users within months.
Overall, the main contributions of our work are as follows:
  • We introduce the Open-domain Urban Itinerary Planning (OUIP) problem, designed to provide personalized urban travel itineraries based on users’ natural language inputs.
  • We develop an LLM-based OUIP system, ItiNera. To the best of our knowledge, this is the first study that combines spatial optimization with LLMs to create spatially coherent, fine-grained urban itineraries tailored to users’ requests.
  • We designed several metrics employing both rule-based assessments and GPT-based metrics to measure the quality and personalization of generated itineraries, which could enlighten future OUIP research by providing a comprehensive metric framework.
  • Extensive experiments on real-world dataset demonstrate that ItiNera delivers the best performance in creating personalized, spatially coherent, high-quality urban itineraries that meet user requirements; ItiNera is also deployed as the core system of TuTu to provide online OUIP service. User feedback validates the effectiveness of our system in real-world scenarios.
Figure 2. An overview of the ItiNera system.
2. Related Works
2.1. LLMs in Urban Application
Since ChatGPT (OpenAI, [n. d.]), Large Language Models (LLMs) have received widespread attention from both the academic and industrial communities. They have demonstrated strong capabilities in tasks such as language processing, instruction-following, planning, and learning. Existing applications of LLMs are concentrated in agent (Wang et al., 2023c; Xi et al., 2023) and prompt (Yao et al., 2023; Wei et al., 2022) design. However, the urban applications using LLMs remain in the early stages. Recent studies highlight the potential of LLMs in urban data processing (Yan et al., 2023), showing their impact on urban planning (Zhou et al., 2024). These works reveal LLMs’ capabilities in predicting human mobility patterns (Mo et al., 2023; Xue et al., 2022), including during public events (Liang et al., 2023), and emphasize their predictive strength (Xue and Salim, 2023; Wang et al., 2023a). In transportation, LLMs contribute to traffic safety analysis (Zheng et al., 2023a), enhance traffic forecasting (de Zarzà et al., 2023), and automate accident report generation (Zheng et al., 2023c), demonstrating their wide applicability in urban transportation. (Zheng et al., 2023b)
Leveraging LLMs for automated travel planning has recently gained public interest. Concurrently, TravelPlanner (Xie et al., 2024) proposes a sandbox environment with various tools for benchmarking LLMs on long-term (minimum 3 days) travel planning, revealing LLMs’ current limitations in handling complex planning tasks. Given this, we present ItiNera, a holistic system designed for OUIP. Unlike TravelPlanner, our system focuses on fine-grained OUIP, addressing urban itinerary planning within a single day, but can be seamlessly extended to multi-day travel planning.
2.2. Itinerary Planning
Current research on IP focuses on creating itineraries based on a set of POIs. Some methods directly optimize the spatial utilities of the itinerary to obtain spatially coherent itineraries, while others define IP as an Orienteering Problem (OP) and consider constraints that include time (Zhang and Tang, 2018; Hsueh and Huang, 2019), space (Rani et al., 2018, 2019), must-see POIs (Taylor et al., 2018), categories (Bolzoni et al., 2014), and their combinations (Gionis et al., 2014; Yochum et al., 2020), optimizing POI utilities such as location popularity (Varakantham and Kumar, 2013; Gunawan et al., 2014) to indirectly ensure the spatial coherence and quality of the itinerary. Although these methods can ensure the quality of POIs in the itinerary, their ability to personalize is limited. Some recommendation-based methods (Ho and Lim, 2022; Tang et al., 2022) could be applied to the IP task. However, these methods often depend on historical user behavior data. Overall, existing IP methods struggle with open-domain, user natural-language inputs, failing to generate personalized itineraries, making them unsuitable for OUIP.
3. Preliminaries
Definition 2 (User Request).
A user request $\small r$ refers to a text string that includes the user’s natural language description of needs for an itinerary. The user request potentially consists of both hard requirements (i.e., constraints) and general preferences for the itinerary.
Definition 3 (Point of Interest).
A Point of Interest (POI) $\small p$ represents a specific location, like landmarks, restaurants, or cultural sites, characterized by its geographical location, category, descriptions, and other relevant context features.
Definition 4 (OUIP-full version).
Given a user request $\small r$ and the user-owned POI set $\small\mathcal{P}=\{p_{j}\}_{j=1}^{N}$, the OUIP problem aims to find an itinerary generator $\small\mathcal{G}_{\Theta}$ to select and order a subset of POIs from $\small\mathcal{P}$ to create a coherent travel itinerary that optimally aligns with the user’s requests $\small r$ while adhering to spatio-temporal considerations:
(1) $\small I\sim\mathcal{G}_{\Theta}\left(\mathcal{P}|r\right),$
where $\small\Theta$ is the parameters of the itinerary generator $\small\mathcal{G}$, and the generated itinerary $\small I$ is an ordered list of POIs.
4. Methodologies
4.1. Overview
In this section, we provide details of ItiNera. As depicted in Fig. 2, we design five modules to enable an OUIP service. We first incorporate a User-owned POI Collection (UPC) module to facilitate travel information gathering and user-owned POI database construction. To plan an itinerary that meets a user’s request, we design a Request Decomposition (RD) module to accurately interpret and extract independent user preferences. Subsequently, we develop a embedding-based Preference-aware POI Retrieval (PPR) Module that fetches the top-K most relevant POIs to form the retrieved POI set for the itinerary. To ensure a spatially coherent itinerary, we employed a Cluster-aware Spatial Optimization (CSO) module to spatially filter and rank retrieved POIs by solving the hierarchical traveling salesman problem, thereby obtaining the candidate POIs. Lastly, the Itinerary Generation (IG) module integrated the ordered candidate POI set with multiple constraints to construct the prompt for LLM’s completion. This approach enables LLM to generate travel routes that are spatio-temporally rational and align with user requests. We provide details for each module in the following subsections.
4.2. User-owned POI Collection
In OUIP, the way of travel information acquisition has undergone a significant transformation from conventional travel planning due to technological advancements. Base on our user and market research, travelers increasingly rely on social media for pre-trip information gathering. Notably, user-generated content is more authentic and current in terms of customer experience than official sources of information. There are existing methods to gather pre-trip information, such as using LLMs to call tools for searching relevant POIs like autoGPT (Gravitas, 2023). However, these LLM-agent-based methods typically consume time and are not suitable for instant use. Moreover, they require users to have relatively specific targets or requirements, which hinder users to discover new POIs.
[…]
Details of this pipeline and the constructed database format are further explained in Appendix E and B, respectively.
4.3. Request Decomposition
Upon receiving user requests, we leverage the advanced natural language processing capabilities of LLMs to extract structured user requests in this module.
Based on our received requests, we’ve observed that a single user request $\small r$ can be decomposed into multiple independent subrequests, each could be generally classified according to their granularity, specificity, and attitude. For granularity, we have (1) POI-level subrequests and (2) itinerary-level subrequests. In terms of specificity, there are (1) specific subrequests and (2) vague subrequests. Regarding attitude, we identify (1) positive subrequests and (2) negative subrequests. Thus, we prompt LLM to decompose the user request $\small r$ base on the above identified categories. We formally use $\small\operatorname{LLM}_{\theta}$ to denote a pre-trained LLM with parameters $\small\theta$:
(2) $\small\mathcal{R}\sim\operatorname{LLM}_{\theta}\left(\,\mathbb{P}_{RD}\left(r\right)\right),$
where $\small\mathbb{P}_{RD}$ wraps the request $\small r$ with instructions and few-shot input-output examples, the result is a set of decomposed independent structured subrequests $\small\mathcal{R}=\{\mathbf{r}_{i}\}_{i=1}^{|\mathcal{R}|}$. Each subrequest $\small\mathbf{r}_{i}$ comprises four keys: ‘pos’, ‘neg’, ‘mustsee’, ‘type’, with their corresponding values $\small\mathbf{r}_{i}^{\operatorname{pos}}$, $\small\mathbf{r}_{i}^{\operatorname{neg}}$, $\small\mathbf{r}_{i}^{\operatorname{must}}$, $\small\mathbf{r}_{i}^{\operatorname{type}}$. Here, ‘pos’ and ‘neg’ are strings indicating the user’s likes and dislikes contained in the subrequest, ‘mustsee’ is a boolean (true or false), and ‘type’ can be one of [”start”, ”end”, ”POI”, ”itinerary”], indicating whether the subrequest targets the starting POI, ending POI, a specific POI, or the entire itinerary, respectively. To explain each field in a subrequest $\small\mathbf{r}_{i}$, we could consider the example illustrated in Fig. 2. The user request, ”I want a historical tour that begins at a cafe, avoiding crowds,” can be broken down into two subrequests. The first specifies the starting point at a cafe, which is a subrequest related to a POI, making it a ”start” ‘type’ with ‘pos’ as ”a cafe”, ‘neg’ as null, and ‘mustsee’ as true, indicating it’s an must visit POI. The second subrequest seeks a historical route that avoids crowds. For this, the ‘type’ is ”itinerary”, ‘pos’ is ”historical tour”, ‘neg’ is ”crowds”, and mustsee is false, indicating a preference rather than a constraint.
For those subrequests identified with the identified ‘type’ as ”POI” and the value of the key ‘mustsee’ is true, we leverage their ‘pos’ strings to conduct fuzzy string match over the POI names stored in the user-owned POI database to retrieve the must-see POIs $\small\mathcal{P}^{\operatorname{must}}=\{p^{\operatorname{must}}_{j}\}_{j=1}^{|\mathcal{P}^{\operatorname{must}}|}$. We provide the full prompt $\small\mathbb{P}_{RD}$ is available in Appendix F.1.
4.4. Preference-aware POI Retrieval
To generate itineraries that meet user requests, the initial step involves selecting appropriate POIs. Although LLMs possess advanced intelligence in language comprehension, their capabilities are limited by their context window size (Liu et al., 2023; Pal et al., 2023), and the cost is proportional to the input tokens.. Given the vast amount of POI data in the real world and the limitations on LLM inference speed that could hinder our system’s response time, we have adopted a embedding-based retrieval approach in this module.
Specifically, consider a single subrequest $\small\mathbf{r}_{i}$, we first employ an embedding model $\small\operatorname{E}_{\theta^{\prime}}$ to encode both the ‘pos’ and ‘neg’ fields:
(3) $\small\mathbf{e}^{\operatorname{pos}}_{i}=\operatorname{E}_{\theta^{\prime}}\left(\mathbf{r}_{i}^{\operatorname{pos}}\right);\,\mathbf{e}^{\operatorname{neg}}_{i}=\operatorname{E}_{\theta^{\prime}}\left(\mathbf{r}_{i}^{\operatorname{neg}}\right),$
where $\small\theta^{\prime}$ denotes the parameters of the embedding model $\small\operatorname{E}$, and $\small\mathbf{e}^{\operatorname{neg}}_{i},\mathbf{e}^{\operatorname{pos}}_{i}\in\mathbb{R}^{d}$ are two resulted $\small d$-dimensional embeddings.
Ideally, we want the queried POIs best fit the positive subrequest while avoiding the negative subrequest. To achieve this, we first use the calculated positive embedding $\small\mathbf{e}^{\operatorname{pos}}_{i}$ to retrieve $\small k$ POIs $\small\mathcal{P}^{\operatorname{pos}}_{i}=\{p^{\operatorname{pos}}_{i,j}\}_{j=1}^{k}$ with top similarity scores $\small\mathcal{S}^{\operatorname{pos}}_{i}=\{s^{\operatorname{pos}}_{i,j}\}_{j=1}^{k}$ and embedding $\small\mathcal{E}^{\operatorname{pos}}_{i}\in\mathbb{R}^{k\times d}$, where $\small s^{\operatorname{pos}}_{i,j}$ and $\small s^{\operatorname{pos}}_{i,j}$ represent the $\small j$th POI and score for $\small i$th positive sub-request, respectively. Then, to ensure avoiding negative POIs, we compute the similarity scores of $\small\mathcal{E}^{\operatorname{pos}}_{i}$ and $\small\mathbf{e}^{\operatorname{neg}}_{i}$ and rerank the retrieved POIs base on the gap between POI-wise positive and negative similarity scores.
Suppose we use $\small\mathcal{E}\in\mathbb{R}^{N\times d}$ to denote the POI embeddings pre-computed over POI attributes that stored in the user-owned embedding database, the above process could be formulate as:
[…]
where the $\small\operatorname{score}(\cdot)$ function measures embedding similarities, and the superscript $\small k$ indicates that the function only returns the top-$\small k$ results based on the similarity scores. The $\small\operatorname{rank}(\cdot)$ reorders POIs based on their similarity scores from highest to lowest.
Lastly, we select the POIs with the highest top-$\small k$ summed-scores from the unioned set of all retrieved POIs to form the final retrieved POI set $\small\mathcal{P}^{rt}$ for the user request $\small r$:
(7) $\small\mathcal{P}^{rt},\mathcal{S}^{rt}=\operatorname{rank}^{k}\left(\cup_{i=1}^{|\mathcal{R}|}\left(\mathcal{P}_{i},\mathcal{S}_{i}\right)\right)\cup\left(\mathcal{P}^{\operatorname{must}},\mathcal{S}^{\operatorname{must}}\right),$
where the set $\small\mathcal{S}^{\operatorname{must}}$ is assigned large values to ensure must-see POIs are included in the final itinerary by subsequent modules. During the union process, the scores of the same POI under different subrequests are summed to form the final score of that POI.
It is worth noting that avoiding negative POIs in the retrieval stage requires scoring across $\small\mathcal{E}^{\operatorname{neg}}_{i}$ and the entire user-owned embedding database, which may hinder the response times. Therefore in actual deployment, we focus on reranking retrieved positive POIs base on negative and positive scores. Given that Eq. 7 ranks all retrieved POIs, our approach does not compromise retrieval effectiveness.
4.5. Cluster-aware Spatial Optimization
A spatially coherent itinerary could essentially enhance the travel experience, travelers tend to move from one cluster of POIs to another (Bolzoni et al., 2014). This behavior can be attributed to the efficiency and convenience of exploring POIs within a proximal area, reducing the time and effort involved in transit. In view of this, spatially filtering all retrieved POIs and sequencing their visitation order is essential in order to inform large language model which so far has exhibited limited capabilities for understanding layout and optimizing routes in multi-dimensional space. To achieve this, we initially compute spatial clusters of the retrieved POIs and then select candidate POIs based on geographical proximity and matching scores. We address cluster-aware spatial optimization by solving a hierarchical traveling salesman problem (Jiang et al., 2014).
4.5.1. Obtain Candidate POI via Spatial Clustering and Filtering
Given the set of retrieved POIs, $\small\mathcal{P}^{rt}$, we create a spatial proximity graph $\small G$ using a distance matrix $\small D$. In this graph, each node denotes a POI, and edges connect nodes (POIs) that are closer than a specific distance threshold $\small\tau$. A community detection algorithm is applied to $\small G$ so that all the nodes are divided into multiple mutually exclusiveclusters. Next, we iteratively select the clusters with the largest summed similarity scores derived from the PPR module until the total number of selected POIs reach a predefined threshold $\small N^{c}$, and the POIs in these clusters form the candidate POIs $\small\mathcal{P}^{c}$ for the user request $\small r$. The above process is detailed in Algo. 1.
Algorithm 1 Spatial Clustering and Selection of Candidate POIs
1:Input: Retrieved POI set $\small\mathcal{P}^{rt}$ with scores $\small\mathcal{S}^{rt}$, Distance threshold $\small\tau$, Candidate POIs num threshold $\small N^{c}$
2:Output: Spatial Clusters $\small\mathcal{C}$, Candidate POIs $\small\mathcal{P}^{c}$
3:// Spatial Clustering
[…]
5:for $\small p_{a}^{rt},p_{b}^{rt}\in V$ with $\small a\neq b$ do
6:  if $\small{distance}(p_{a}^{rt},p_{b}^{rt})<\tau$ then
7:   $\small E\leftarrow E\cup\{(p_{a}^{rt},p_{b}^{rt})\}$
8:  end if
9:end for
10:while $\small V\neq\emptyset$ do
11:  $\small c\leftarrow\text{largest clique in }G$
12:  $\small\mathcal{C}\leftarrow\mathcal{C}\cup\{c\}$; $\small V\leftarrow V\setminus c$
13:end while
14:// Selection of Candidate POIs
15:for each cluster $\small c\in\mathcal{C}$ do
16:  $\small s^{rt}_{c}\leftarrow\sum_{p_{j}\in c}s_{j}^{rt}$
17:end for
18:Sort $\small\mathcal{C}$ in descending order of $\small\mathcal{S}^{c}=\{s^{rt}_{c}\}_{c=1}^{\mathcal{C}}$
19:while $\small|\mathcal{P}^{c}|<N^{c}$ do
20:  $\small c_{\text{max}}\leftarrow\arg\max_{c\in\mathcal{C}}s_{c}^{rt}$
21:  $\small\mathcal{P}^{c}\leftarrow\mathcal{P}^{c}\cup c_{\text{max}}$; $\small\mathcal{C}\leftarrow\mathcal{C}\setminus\{c_{\text{max}}\}$
22:end while
23:return $\small\mathcal{C},\mathcal{P}^{c}$
4.5.2. POI Ordering via Solving a Hierarchical TSP
After obtaining the spatial clusters $\small\mathcal{C}$, we further optimize the visitation order of the candidate POIs to ensure a spatially coherent itinerary. This process first involves determining the access order of each cluster. Then within each cluster, we determine the visitation order of POIs by solving TSPs with starting and ending POI constraints. Specifically, the start and end points for each cluster are selected based on the proximity of POIs in adjacent clusters as described in Fig. 2. We provide an Algo. 2 below that illustrates the procedure.
[…]
Algorithm 2 Hierarchical TSP for POI Ordering
1:Input: Spatial clusters $\small\mathcal{C}$, candidate POIs $\small\mathcal{P}^{c}$, distance matrix $\small D$
2:Output: Ordered list of candidate POIs $\small\mathcal{P}^{\text{order}}$
3:// POI Ordering
[…]
5:for each cluster $\small c$ in $\small\mathcal{C}^{\text{order}}$ do
6:  $\small p_{\text{start}}^{c},p_{\text{end}}^{c}\leftarrow\text{GetClusterEndpoints}(c,\mathcal{C}^{\text{order}},D)$
7:  $\small c_{\text{path}}\leftarrow\text{SolveConstraintTSP}(c,p_{\text{start}}^{c},p_{\text{end}}^{c},D)$
8:  $\small\mathcal{P}^{\text{order}}\leftarrow\mathcal{P}^{\text{order}}\cup c_{\text{path}}$
9:end for
10:// Start POI Selection and POI Reordering
[…]
13:return $\small\mathcal{P}^{\text{order}}$
4.6. Itinerary Generation
Selecting any subset from $\small\mathcal{P}^{\text{order}}$ guarantees a spatially coherent itinerary, but a high-quality itinerary should also comply with additional constraints, such as matching the user’s time availability and ensuring practicality. For example, it should avoid consecutive restaurant visits and ensure venues are visited at appropriate times, such as not planning bars in the morning or coffee shops late at night. Given the complex nature of these constraints, traditional optimization-based algorithms can become excessively complicated and lack variability (Yochum et al., 2020; Taylor et al., 2018), hindering system deployment and itinerary diversity. To address this, considering the advanced reasoning and planning capabilities of LLMs, we leverage LLMs to generate final itineraries that satisfy these diverse constraints.
Simplified prompt $\small\mathbb{P}_{IG}$ for final itinerary generation. See Appendix F.4 for full version. Please consider carefully and use the provided Candidate POIs list to craft a one-day itinerary in the form of an engaging and realistic travel story. ## Itinerary Information: - **Candidate POIs**: {p_order} 1: Time Cafe, Dongcheng District. Located in the bustling Nanluoguxiang, this cafe is a popular spot for photo enthusiasts and also a cat cafe. Surrounded by beautiful roses, its perfect for capturing memories.“ 2: - **Must-see POIs**: {r_must} - **Keyword Requirements**: {r_pos} - **Users Original Request**: {r} - **Start POI**: {p_start} - **End POI**: {p_end} ## Constraints: - Only choose POI from the **Candidate POI** list and in ascending order. - Bars should be at the end of the itinerary, while coffee shops should not be the last POI. ## Task Description: 1. You have approximately {hours} hours. Select suitable POIs from the **Candidate POIs** list in ascending order, ensuring your itinerary is selectively filtered and does not include all POIs from the list. 2. All POIs added to the itinerary must follow the ascending order of **Candidate POIs**. 3. Generate a JSON file containing all selected POIs. ## Output Format: {{ itinerary”: List of POIs, separated by ’->’” Overall reason”: Overall recommendation reason for the designed day trip itinerary”, pois”: {{ n”: Description and recommendation reason for each POI”, }} }}
Specifically, the primary objective of this module is to effectively utilize LLM to select an optimal subset from $\small\mathcal{P}^{\text{order}}$, which closely aligns with user requests while adhering to various constraints. This process can be formally defined as follows:
[…]
Utilizing an LLM to generate the final travel itinerary allows for a balance among various considerations. Additionally, leveraging the understanding and reasoning capabilities of LLM, we can create a time-appropriate travel itinerary tailored to the user requests. Specifically, we leverage a simple time indication prompt $\small\mathbb{P}_{TI}$ to indicate the travel time (in hours) of the itinerary, and the results can be integrated into the “Task Description” part of the prompt $\small\mathbb{P}_{IG}$. We provide the complete $\small\mathbb{P}_{TI}$ and further details on the travel time considerations of the itinerary in Appendix F.2.
5. Experiments
To evaluate the performance of ItiNera for OUIP, we conduct extensive experiments to answer the following research questions: RQ1, why does directly using LLM to generate travel itineraries not guarantee a good user experience, and what advantages can ItiNera offer in comparison? RQ2, how do the various components of ItiNera enhance the quality of the generated itineraries? RQ3, how well does ItiNera serve the OUIP service at TuTu in an online production environment? Sec. 5.3 provides results to answer RQ1. Sec. 5.4 provides an ablation study to answer RQ2. Sec. 5.6 provides online performance results to answer RQ3.
5.1. Experimental Setup
5.1.1. Data Description
Shanghai Qingdao
[…]
PQ IQ Match
[…]
PQ IQ Match
IP (Gunawan et al., 2014) 6.4 1573.3 2.96 / 30.3 26.2 17.8 7.6 4134.3 2.86 / 23.6 16.8 20.2
Ernie-Bot 4.0 (Sun et al., 2021) 15.7 513.5 0.91 15.2 42.1 46.5 42.5 27.2 776.2 0.78 21.4 43.4 38.2 33.3
GPT-3.5 (OpenAI, [n. d.]) 16.6 422.4 0.83 13.5 40.4 43.1 45.4 25.5 691.5 0.55 22.0 33.4 39.0 46.6
GPT-4 (OpenAI, 2023) 18.0 267.2 0.56 8.2 45.0 48.2 46.9 27.3 569.4 0.49 19.6 46.6 48.7 48.4
GPT-4 CoT (Wei et al., 2022) 18.4 258.3 0.49 7.5 / / / 30.2 542.6 0.43 17.8 / / /
Ours 31.4 86.0 0.42 / 69.8 64.6 72.0 35.4 225.8 0.26 / 71.2 68.2 67.8
Beijing Hangzhou
[…]
PQ IQ Match
[…]
PQ IQ Match
IP (Gunawan et al., 2014) 3.3 3034.2 2.26 / 27.8 18.2 20.4 1.8 1744.4 1.52 / 34.8 31.4 22.5
Ernie-Bot 4.0 (Sun et al., 2021) 18.8 379.4 0.74 12.8 31.2 34.8 32.1 12.9 605.2 1.17 24.4 43.6 34.3 38.2
GPT-3.5 (OpenAI, [n. d.]) 19.7 347.8 0.58 14.3 29.2 40.5 43.8 14.4 665.4 1.16 19.8 41.2 40.8 32.8
GPT-4 (OpenAI, 2023) 20.6 342.6 0.52 11.1 45.4 43.6 45.2 14.8 746.1 1.09 23.2 46.2 39.6 39.4
GPT-4 CoT (Wei et al., 2022) 21.0 327.7 0.54 10.2 / / / 15.5 455.0 1.09 18.6 / / /
Ours 28.4 79.2 0.46 / 59.2 67.6 75.2 21.4 30.5 0.12 / 61.6 65.4 68.3
Table 1. Overall results. GPT-based metrics represent the win rate against GPT-4 CoT.
We conduct experiments on an urban travel itinerary dataset from four Chinese cities, collected in collaboration with a leading travel agency specializing in single-day city tours. The dataset comprises top-rated 1233 urban itineraries and 7578 POIs. Each data sample contains a user request, the corresponding urban itinerary plan, and detailed POI data. For an in-depth understanding of the data format, sample entries, and key preprocessing methodologies employed, refer to Appendix B.
5.1.2. Implementation Details
We utilize two LLMs through their APIs: GPT-3.5 (gpt-3.5-turbo version) and GPT-4 (gpt-4-turbo version). GPT-4 is used for the final itinerary generation ($\small\mathbb{P}_{IG}$) to guarantee the overall quality of the itinerary, and GPT-3.5 is employed for other LLM interactions to ensure the response speed of our OUIP service. We also incorporate the text-embedding-ada-002 model for embedding purposes. The spatial coherence of itineraries is optimized through an open-source TSP solver [^1] https://github.com/fillipe-gsm/python-tsp. The integration of POI data encompasses various attributes, including geographical coordinates, user ratings, categorizations, and physical addresses, facilitated through the Amap API [^2] https://lbs.amap.com/. For the production environment, our infrastructure leverages Amazon Elastic Container Service (ECS) to orchestrate a cluster of servers, each provisioned with an 8-core CPU and 16GB of RAM, optimizing performance for Python-based applications and machine learning APIs. Notably, the original languages of our service and data are simplified Chinese, and the translated version is used for demonstration purposes in this paper. More implementation details can be found in Appendix C.
[…]
Ours w/ GPT-3.5 27.6 79.4 0.56 67.3 58.8 61.4
Ours (full) 31.4 86.0 0.42 69.8 64.6 72.0
Table 2. Ablation study on Shanghai dataset.
5.2. Evaluation Metrics
According to the concept of the OUIP task, a satisfactory itinerary must possess the following characteristics: spatial coherence and alignment with the user’s needs. To this end, we have designed several objective metrics to evaluate the generated itineraries.
5.2.1. Rule-based Metrics
We first design some rule-based metrics: (1) Recall Rate (RR): the recall rate of POIs in the ground truth itinerary; (2) Average Margin (AM): The average difference per POI between the total distance of the generated itinerary and the shortest distance (calculated by TSP); (3) Overlaps (OL): the number of self-intersection points in the generated itinerary, e.g. there are 2 overlaps in Fig. 3 (right). (4) Fail Rate (FR): The percentage of POIs generated by LLM that cannot be matched with the queried POIs from the map service (serve as the full POI set). RR evaluates the method’s accuracy in understanding user requests and recommending correct POIs. AM and OL measure spatial optimization for POI visit order, where lower is better. FR evaluates LLM’s information accuracy and indicates the method’s compatibility with the online service. Locations of failed POIs are inaccessible and cannot be displayed on the map, which lowers the user experience.
5.2.2. GPT-based Metrics
The above rule-based metrics are intuitive. Some other aspects are hard to quantify, e.g. the intrinsic appeal of the POIs or the degree of alignment between user requests and the generated itinerary. Therefore, following previous benchmarks for automatic evaluation (Li et al., 2023; Bang et al., 2023; Wang et al., 2023b), we propose several GPT-based metrics: (1) POI Quality (PQ): how interesting and diverse the POIs are; (2) Itinerary Quality (IQ): the overall quality and coherence of the itinerary; (3) Matchness (Match): the matchness between the itinerary and the user request; We ask GPT-4 to rank two candidate itineraries generated with different methods and compute the win rate. We repeat the same query for at least 10 times with different seeds to obtain a reliable result.
Automated evaluation through GPT offers a more efficient and cost-effective alternative to human evaluation for large-scale experiments. For example, the cost of GPT tokens for the entire ablation study in Sec. 5.4 is less than \$25. In addition, the results of the online subjective evaluation in real-world scenarios in Sec. 5.6.2 demonstrate consistent results between human and GPT assessments.
5.3. Overall Results
5.3.1. Baselines
We compare ItiNera with the following baselines:
  • IP (Gunawan et al., 2014): A traditional IP method. We simplified it to use LLM for time budgeting and to consider POI ratings as utilities.
  • Ernie-Bot 4.0 (Sun et al., 2021): The best-performing model on Chinese LLM tasks, selected as our dataset and service are in Chinese.
  • GPT-3.5 (OpenAI, [n. d.]): The standard ChatGPT model.
  • GPT-4 (OpenAI, 2023): The enhanced model offering superior performance and broader knowledge coverage.
  • GPT-4 CoT (Wei et al., 2022): The model improves GPT-4 with the LLM-as-agent method, Chain-of-Thought reasoning.
We employ the same itinearry generation prompt for all baselines, including basic task requirements and output format as in $\small\mathbb{P}_{IG}$. Notably, for GPT-4 CoT, we extend the prompt by integrating the “thoughts”. These prompts are detailed in Appendix F.5. More details on baseline implementations can be found in Appendix C. Note that the baseline IP and our method does not compute Fail Rate as the candidate POIs are from the dataset.
5.3.2. Results Analysis
We provide analysis for results (RQ1). As illustrated in Tab. 1, our proposed ItiNera significantly outperforms all baselines across all metrics. For rule-based metrics, ItiNera shows ~30% improvement over the best baseline, highlighting its superior capability in personalizing user experiences. Furthermore, it maintains spatial coherence by generating itineraries that are only marginally longer (~100 meters per planned POI) than the shortest path solved by TSP, a remarkable achievement considering potential user requests that may deviate from spatially optimal routes. Notably, ItiNera is the only method to surpass the performance of GPT-4 CoT in GPT-based evaluations, particularly excelling in matchness. These findings underscore ItiNera’s efficacy in enhancing spatial coherence and aligning with user requests in OUIP.
5.4. Ablation Study
To understand how the different components of ItiNera affect the quality of the generated itineraries (RQ2), we compare with the following variants of ItiNera:
  • GPT-4 CoT + UPC: This variant integrates the UPC module to LLMs to generate itineraries conditioned on user-owned POIs.
  • Ours w/o RD: This variant leverages the entire user input string’s embedding to retrieve the POIs.
  • Ours w/o PPR: This variant quantifies the contribution of the PPR module compared to our full system.
  • Ours w/o CSO: This variant removes the CSO module and let the LLM called in the IG module determines the order of candidate POIs when generating the final itineraries.
  • Ours w/ GPT-3.5: This variant replace the GPT-4 with GPT-3.5 for generating the final itinerary.
We remove Fail Rate in ablation study, since all variants equipped with UPC never generate POIs not presented in the database.
The results are shown in Tab. 2. Comparing the first two rows, we find UPC can improve Recall Rate and Matchness of the GPT-4 CoT baseline. The variants “w/o RD”, “w/o PPR”, and “w/ GPT-3.5” exhibit lower Recall Rate, POI Quality and Matchness than our full model. This is because they are relatively inept at understanding user requests and selecting relevant POIs. However, they display lower Average Margin and Overlaps. We attribute this to the idea that there is a trade-off between spatial optimization and matchness. Imposing constraints to align a system with human preference will sacrifice its inherent ability. This is similar to conditional generation tasks in other domains (Saharia et al., 2022; Di et al., 2021) , where aligning output with human input hinders the generation quality and is worth than unconditional generation. This trade-off is also observed when removing CSO module (“w/o CSO”) leads to worse Average Margin and Overlaps but better Recall Rate, POI Quality and Matchness. Our full model manages to balance matchness with spatial ability.
In addition to effectiveness of the PPR module as verified by the results of “w/o PR”, the integration of the PPR module could also help us address the issue of the limited sizes of LLMs’ context windows, saving costs on LLM for our system. Lastly, we find that the results of the variant “w/ GPT3.5”, although compromised, still significantly outperform the GPT-3.5 baseline in Tab. 1, demonstrating our system’s adaptability to alternative LLMs for OUIP.
5.5. Qualitative Results
We further conduct a qualitative study to demonstrate the effectiveness of our system and the importance of integrate LLM with a CSO module. Consider a user request “I’m seeking an artsy itinerary that includes exploring the river’s bridges and ferries.” We visualize the results from ItiNera and GPT-4 CoT in Fig. 3.
Figure 3. Itineraries of ItiNera (left) and GPT-4 CoT (right).
We find that our itinerary better matches the user preferences. The itinerary passes several bridges along the Huangpu River, include a ferry crossing, and conclude at the art-atmosphere-rich Duoyun Bookstore, offering a restful endpoint for users. In contrast, the POIs selected by GPT are more mainstream. Moreover, our spatial arrangement is more logical, avoiding detours and concentrating selected POIs within two spatial clusters. The itinerary generated by GPT is spatially poor, with a disordered sequence of visits and contains excessively distant POIs. Beyond this example, GPT also risks hallucinating non-existent POIs, highlighting the superiority of our system ItiNera in comparison.
5.6. Online Performances
5.6.1. Online Deployment
Our system has been deployed as the core algorithm of the TuTu online OUIP service. Within a month of operation, it has attracted over 1K registered users who have created more than a thousand itineraries across 60 cities, incorporating over 30K user-owned POIs. We provide demonstration of the deployed service in Appendix A.
In our online production, the itinerary generation utilizes the GPT-4 API with streaming. The average latency for the initial response token is around 6 seconds, and the complete response takes around 20 seconds. Our designed prompt asks for a short output, thereby reducing the overall response time.
5.6.2. Online Evaluation
Method Rule-based POI Quality Itinerary Quality Matchness
AM OL Expert User GPT Expert User GPT Expert User GPT
GPT-4 CoT 511.4 0.79 3.2 3.6 30% 2.5 3.0 32% 2.9 2.6 28%
Ours 107.6 0.44 3.8 4.3 70% 3.2 3.8 68% 3.6 3.5 72%
Table 3. Online evaluation.
To verify the effectiveness of our system in real-world scenarios, we conduct online evaluations (RQ3). Subjective evaluation via user studies has been extensively employed in prior research on generative tasks (Saharia et al., 2022; Rombach et al., 2022; Zhuo et al., 2023) where objective metrics fail to adequately assess specific dimensions of the output quality. The TuTu online OUIP service provides a natural way to conduct subjective evaluation. For a small group of users (User), we provide them with a feedback interface upon the completion of each itinerary generation. For those users, we randomly replace the algorithm in the service with the baseline GPT-4 CoT to generate itineraries in the probability of 50%. In the same way as the GPT-based metrics, users are required to rate the POI Quality, Itinerary Quality and Matchness, following the Likert scale (1 for poor and 5 for excellent). We gathered 1443 results from 464 users. These ratings are subsequently utilized to upgrade our system to enhance user experience. We also invited 33 experienced travel assistants (Expert) from our collaborative travel agency to provide evaluation on the collected 1443 queries and itineraries from a professional perspective. For GPT-based metrics, we first use the other algorithm to generate an itinerary for the same query to compare with the existing itinerary. Then, we ask GPT-4 to rank the two candidate itineraries and compute the win rate over all queries.
The average evaluation results are reported in Tab. 3. One can observe that our method is preferred by experts and online users across all metrics, especially for Matchness, validating the effectiveness of our system in real-world scenarios. Moreover, the online evaluation results are consistent with the GPT evaluation win rate, indicating that the proposed GPT-based metrics are appropriate and adaptable where massive results need to be evaluated but rule-based evaluation is not sufficient.
6. Conclusion
We introduce the problem of OUIP and a novel solution approach ItiNera that integrates Large Language Models (LLMs) with spatial optimization. ItiNera enables the generation of personalized and spatially coherent itineraries directly from natural language requests. Our system is deployed in the TuTu online service. The results from offline and online experiments validate the effectiveness of our approach. This study not only establishes a new benchmark for itinerary planning technologies but also broadens venues for further innovations in leveraging LLMs for complex problem-solving in urban contexts. Future studies will focus on categorization of diverse user requests, and also further enhance the system’s scalability and real-time performance, to better serve the growing user base of the TuTu online service.
Acknowledgement
We thank Zhaofeng Wu and Han Zheng at MIT, and Tiange Luo at UMich, for participating in discussions about the initial idea and providing valuable insights on this project.
Appendix A Demonstration of the Deployed Service
Figure 4. Screenshots of the deployed service: POI view & Itinerary view.
We provide screenshots of our deployed system in TuTu online service in Fig. 4. The left screenshot shows the POI interface, where users can add new POIs by direct searching or pasting a link of a travel-related post. They can filter their desired POIs to display on their personal map. The POI icon represent its category (entertainment, restaurant, etc.). By pressing the bottom right button, users can select several POIs to create an itinerary. They can also use our proposed system to generate an itinerary from natural language input (the left figure of Fig. 1).
The right screenshot shows the itinerary interface. Users can browse the itineraries they have created and generated. They can tap one itinerary see the details (the middle and right figures of Fig. 1).
Appendix B Dataset
In this section, we provide the data format of the collected real-world dataset. Specifically, the data for each city contains two tables: one is the POI table, which primarily stores the POIs and their features, and the other is the Itinerary table, which is used to store users’ natural language requests and the corresponding ground truth itineraries.
id name address city description longitude latitude rating category context
1 The Bund Zhongshan East 1st Rd, Huangpu Shanghai The Bund is a waterfront area … 121.4906033011 31.2377704249 5.0 site
Table 4. POI data sample.
[…]
user_request itinerary
I’m seeking an artsy itinerary that includes exploring the river’s bridges and ferries. [1, 3, 6, …]
Table 5. Itinerary data sample.
The sample itinerary data is shown in Table 5, which contains only two columns: one for the user’s request and the other storing a list of POI IDs that represent the ground truth itinerary (label) for the user’s request.
Appendix C Implementation Details
C.1. Baseline Settings
For the baseline IP, we adopt the $\small\mathbb{P}_{IT}$ for time budgeting. For the LLM baselines, we prompt them to generate itineraries based on user requests. For each POI in the generated itinerary, we search for it using the Map API. Here, the database associated with the Map API is considered to be the current collection of all existing POIs. We leverage fuzzy string matching  [^3] https://github.com/seatgeek/thefuzz to determine if there is a match with specific POIs. The failed POIs contribute to the failure rate metrics. For the matched POIs, attributes of the POI (such as location) are attached to the itinerary for subsequent evaluation.
C.2. Production Environment Details
In our production environment, we leverage the AWS ECS service to orchestrate a cluster of servers, each provisioned with an 8-core CPU and 16GB of RAM, optimizing the performance for Python-based applications and machine learning APIs. To ensure high availability and enhanced data input/output operations, we employ a multi-region Amazon RDS setup, complemented by Redis  [^4] https://redis.io/ for advanced caching strategies. Additionally, we integrate Amazon OpenSearch Service  [^5] https://aws.amazon.com/opensearch-service/ for executing complex queries and supporting sophisticated search and embedding functionalities. To manage user AI API requests efficiently, we’ve implemented a message queue-like system using Amazon Simple Queue Service (SQS), which offers a highly scalable, reliable, and secure means of message queuing and inter-process communication.
Appendix D Cluster-aware Spatial Optimization Supplementary
We present the implementation details of algorithms involved in cluster-aware spatial optimization.
D.1. SolveTSP
‘SolveTSP’ implements simulated annealing algorithm for efficiently solving the TSP problem with a large set of candidates. Simulated annealing is a classic metaheuristic approach where the model iteratively proposes new solution and replaces the current solution if a certain condition is satisfied until the temperature goes down to zero. We detail the implementation for simulated annealing in Algo 3.
Algorithm 3 Simulated Annealing for TSP
1:procedure SimulatedAnnealing(cities, $\small T_{init}$, $\small T_{min}$, $\small\alpha$)
2:  $\small solution\leftarrow\text{RandomSolution}(cities)$
3:  $\small T\leftarrow T_{init}$
4:  while $\small T>T_{min}$ do
5:   $\small newSolution\leftarrow\text{Neighbor}(solution)$
6:   $\small costDifference\leftarrow\text{Cost}(newSolution)-\text{Cost}(solution)$
7:   if $\small costDifference<0\text{ or }\exp(-costDifference/T)>\text{Random}()$ then
8:     $\small solution\leftarrow newSolution$
9:   end if
10:   $\small T\leftarrow\alpha\times T$
11:  end while
12:  return $\small solution$
13:end procedure
  • RandomSolution: Generates a random permutation of the cities as the initial solution.
  • Neighbor: Produces a new solution by making a small change to the current solution. In our implementation, we consider four types of operations including swapping two randomly selected cities, inverting a subroute, inserting a randomly selected city to another position, and inserting a randomly selected subroute to another position.
  • Cost: Calculates the total distance of the proposed solution’s path.
  • $\small T_{init}$ and $\small T_{min}$: The initial and minimum temperatures for the SA algorithm. In our implementation, $\small T_{init}$ is set to 5000 and $\small{T_{min}}$ is set to 0.
  • $\small\alpha$: The cooling rate that determines how fast the temperature decreases. In our implementation, $\small\alpha$ is set to 0.99.
D.2. SolveTSPWithEndpoints
Algorithm 4 SolveTSPWithEndpoints
[…]
2:Solve the following linear program:
Minimize: $\small\displaystyle\min\sum_{i\neq j}x_{ij}\cdot dist[i][j]$
[…]
Subject to: $\small\displaystyle\sum_{i\neq k}x_{ik}=1,\quad\forall k\neq s,e$
$\small\displaystyle\sum_{i\neq k}x_{ki}=1,\quad\forall k\neq s,e$
$\small\displaystyle{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}//\text{Add constraints for source node and sink node}}$
$\small\displaystyle\sum_{i\neq s}x_{si}=1$
$\small\displaystyle\sum_{i\neq s}x_{is}=0$
$\small\displaystyle\sum_{i\neq e}x_{ie}=1$
$\small\displaystyle\sum_{i\neq e}x_{ei}=0$
$\small\displaystyle{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}//\text{Eliminates all subtours}}$
$\small\displaystyle\sum_{i\in S}\sum_{j\notin S,j\neq i}x_{ij}\leq|S|-1,\quad\forall S\subset\{1,\ldots,n\},S\neq\emptyset,S\neq\{1,\ldots,n\}$
$\small\displaystyle{\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}//\text{Add binary variable constraints}}$
$\small\displaystyle x_{ij}\in\{0,1\},\quad\forall i,j$
3:Return Optimal path
In each cluster, the dataset typically comprises a limited set of candidate points. Consequently, the prioritization shifts towards optimizing the accuracy of the resultant solution, rather than focusing solely on the computational efficiency. To address the Traveling Salesman Problem (TSP) with predetermined starting and ending points, we adopt a linear programming (LP) methodology. We detail the formulation of linear program in Alg. 4.
Appendix E Overview of the POI extraction pipeline
As illustrated in Sec. 4.2, we design a pipeline to automatically extract POIs and relevant information from user-generated content on various social media platforms. The pipeline consists of the following steps:
  • Scrape text, images, and videos from the input link of a travel-related post.
  • Use Automatic Speech Recognition (ASR) to obtain transcription from the video, and Optical Character Recognition (OCR) to extract text from the images. Merge them with the original text to obtain the post information.
  • Use GPT-3.5 to extract POI names and locations from the post information.
  • Use map service API to look up the extracted POI names and obtain the coordinates and extract POI names, similar to the evaluation pipeline in Sec.5.1.2.
  • Use GPT-3.5 to generate POI descriptions from the POI names and post information.
The prompt for extracting POI names and locations is provided below.
Prompt for Extracting POI Names and Locations # Guidelines ## Task Background Your task is to identify and extract mentioned cations in the posts/travelogues provided by users to help them quickly find these places on a map. Now, based on the content of the post and in context, carry out the extraction and description of Points of Interest (POIs) mentioned in the post. Focus primarily on places that can be visited, rather than merely on place names. ## Notes on Handling POIs 1. Comprehensive Definition of POI: Typically used to describe a specific geographical location or site, such as restaurants, hotels, streets, attractions, museums, bars, cafes, malls, etc. These locations or sites may have specific value or interest to users or travelers. 2. Characteristics of POI: Specific places recommended or mentioned in the post that are usable for dining, entertainment, etc. 3. Specificity: A POI refers to a specific, particular place, not a broad geographical area or city name. 4. Uniqueness: When a text is separated by symbols like ”/”, ”&”, ”,”, for example, Julu Road/Tianzifang”, it often represents two POIs, in this case, Julu Road and Tianzifang should be extracted separately. 5. Examples of POI: Specific restaurants, performance venues, attractions, shops, streets, etc. 6. Non-POI Examples: Collections of places, food names, types of cuisine, performance groups, exhibition events, etc. ## Post Structure Title: The posts title. Text: The main body content of the post. Text in the images: text recognized from the images. Transcribed text: text transcribed from the video. ## Task Process 1. Extraction: Based on your reasoning, judgment, and knowledge, extract all mentioned POIs from the post. 2. Verification: In the context of each POI, ensure all POIs fit the definition and are specific places. 3. Address Information: In the context of each POI, find related address information that can be searched on a map, such as ”158 Julu Road, Shanghai.” 4. Handling No Information: If no location information is available, return an empty POI list: {{}}. 5. Formatting: Organize information into the specified JSON structure. ## Output Format ### Specific Format {{ POI Name”: Related Address Information for the POI }} ### Examples Example 1: If the original post mentions Lao Nong Tang Noodle Shop in Luxi: A time-honored noodle shop that appears on Shanghais must-eat list all year round!”, the output for this POI should be {{ Lao Nong Tang Noodle Shop in Luxi”: null }} Example 2: If the original post mentions Red Baron (Jianye District Wentiyi Road branch) Looking around, the most striking red on the entire Wentiyi Road, seamlessly blending with Mixue Bingcheng”, the output for this POI should be {{ Red Baron (Jianye District Wentiyi Road branch)”: null }} ## Output Standards - The output is a dictionary, with keys being the POI names and values being the related address information for the POI. If address information is missing, please use null to fill in. - Ensure the output is in valid JSON format and can be parsed by Python json.loads. ## Task Start Please begin processing the post content: “‘{post_info}“‘. Note: Ensure the return format follows {{Point of Interest Name: Related Address Information}}. Ensure it can be json.loads parsed.
The prompt for generating POI descriptions is provided below.
Prompt for Generating POI Descriptions {post_info} Based on the content of the above post, please write out the reasons for recommending each location to tourists in the following list. Consider what can be done at this location, its features, and why it is fun. If the original post lacks information, you may appropriately supplement based on your knowledge, but please ensure brevity. The related information for each point should not exceed 30 words. The results should be output in JSON format, specifically in the form {{Place Name: Information related to the place from the original post}}, where Information related to the place from the original post should be a sentence or phrase, like a string. If a place does not have any relevant information, fill in the description related to the place from the original post with null”.
We execute an automated process to extract POIs from the most recent trending posts and update a comprehensive POl database. At a regular interval of 24 hours, we obtain recent trending travel-related posts across multiple cities on social media platforms, and run the above pipeline to extract POI names, locations and descriptions to maintain the database.
Appendix F Prompts
In this section, we provide all the prompts for LLMs used in our method and evaluation.
F.1. Prompt for Decomposing User Requests
Prompt for Decomposing user requests Please help me break down a user request into multiple independent requirements, each including both positive and negative requirements. Return the results in the following format directly based on the **User Request** given, without writing any code. ### Output Format: Return a list, where each item is a dictionary representing an independent requirement, with the following key-value pairs: - **pos**: The positive requirement, representing what the user wants, excluding any negative requirements. - **neg**: The negative requirement, generally what the user does not want, dislikes, or refuses. All negative requirements must be captured in this field, for example, non-spicy should extract spicy”, dont want crowded places should extract crowded”, hate noisy should extract noisy”. - **mustsee**: Indicates whether this requirement represents a specific place name. If so, this field is true‘, otherwise, it is false‘. - **type**: Indicates whether the requirement is for a place or an itinerary”, with place having sub-types place”, starting point”, and ending point”. Overall, this field can have the values place”, starting point”, ending point”, or itinerary”. - Your return should be a list in the following format: [ {{ pos”: positive requirement”, (excluding negative requirements) neg”: negative requirement (whats not wanted, disliked, refused, not wanting to go or see, any negated requirement), mustsee”: true (whether its a must-see point, all specific places should be set to true), type”: place }}, ] - The **positive requirement** must not be empty, and it must not include any negative requirements. All negative requirements should be summarized in the value of the neg field. - Set to null in cases where there are no **negative requirements** for a specific place. - Sometimes users only describe what they do not want (negative requirements), in such cases, you should summarize a **positive requirement** based on the **negative requirement**. For example, if a user says dont want spicy food’, the output should include: pos corresponding to food”, neg corresponding to spicy”. - Independent requirements must have specific descriptions or demands to be considered a requirement, for example, recommend a route does not count as an independent requirement. - mustsee must be a specific place name, not a general term. - If a place is definitively a starting point or ending point”, then the value of the type field should be starting point or ending point”, respectively. Starting points and ending points are must-see points, with the mustsee field set to true. - A place can only be considered as a starting point or ending point if it is a specific attraction or location, and there can only be at most one starting point and one ending point”. - The return should not include any other content. ### Example Outputs: Example 1: User Request: I want to start by visiting Sinan Mansions, then find something fun to do nearby, and I dont want it to be crowded Output: [ {{ pos”: Sinan Mansions”, neg”: null, mustsee”: true, type”: starting point }}, {{ pos”: fun places near Sinan Mansions”, neg”: crowded”, mustsee”: false, type”: place }} ] ### mustsee Field Assignment Examples mustsee true for specific place names: Hualian Supermarket”, Old Mac Coffee Shop”, Wukang Mansion”, Nanluoguxiang”, mustsee false for general place names: supermarket”, milk tea shop”, bar”, coffee”, ### Output Guidelines - Return a list, each item in the list is a dictionary containing pos”, neg”, mustsee”, and type key-value pairs. - Return as a JSON List. - The list can be empty; if empty, just return a JSON list. - The output should not include any other information, ensuring it can be parsed by json.loads. ### User Request {user_req} ### Task Overview Your task is to analyze and break down the **User Request** into independent requirements and return them. 1. First, separate the different independent requirements, breaking down each into positive and negative requirements. 2. Positive requirements should only include what the user wants, and negative requirements should only include what the user does not want. 3. For each independent requirement, refer to **mustsee field assignment examples** to assign a value to the mustsee field, analyzing whether the **positive requirement** is a specific place name. If so, set mustsee to true, otherwise set it to false. 4. Refer to the **examples** and **output format** to complete the other fields. #### Notes: - Do not include duplicate independent requirements; ensure each independent requirement corresponds to different key points in the users needs. - Itinerary requirements should be for the whole itinerary, such as including several places, approximate time, etc., all others are place requirements. - The type field can only be one of [”place”, itinerary”, starting point”, ending point”]. - All attractions must be completely separated, such as Nanluoguxiang and Drum Tower must be split into Nanluoguxiang and Drum Tower as two requirements. Now, based on the **user Request**, refer to the **example outputs**, and return according to the **output guidelines** and **output format**.
F.2. Prompt for Indicating Travel Time of an Itinerary
Prompt for Travel Time Indication Please play the role of a top AI Travel Time Planning Assistant. Your job is to determine the time needed for a days itinerary based on the user request. If the user request is empty, please default to [”4”]. ## Task Overview Your task is to return the required time for a days itinerary based on given the user request. If the user request mention specific time constraints for the route, return the itinerary time directly as per the users request, up to a maximum of 8 hours (return [”8”] if it exceeds 8 hours). Please return the itinerary time directly based on the user request, no need to write any code. ## User Request {user_reqs} ## Input-Output Examples - **Example 1**: - **User Request**: Id like to visit a museum, enjoy authentic cuisine, and experience nightlife.” - **Output**: [”8”] - **Example 2**: - **User Request**: I want to tour historical buildings and take in the city views - **Output**: [”6”] - **Example 3** (In this example, the user specifies approximately **five hours** for the route): - **User Request**: I plan to explore the Huangpu River and Yu Garden for about five hours.” - **Output**: [”5”] ## Output Specifications - Return a list of length 1, containing a single integer representing the required itinerary duration (in hours). The value range is 1 to 8. - Return as a JSON List with only one element inside. - The list can be empty; please only return a JSON list. - Ensure your output contains no additional information and can be parsed by json.loads. Now, based on the **User Request**, return the time required for a days itinerary according to the **Output Specifications**.
In this work, we utilize the inference capability of LLMs to estimate the duration of an itinerary based on a user request, which is used to instruct the IG module to generate an itinerary with a reasonable duration. For more complicated considerations, such as stay duration and travel time between POIs, we leave them for future research.
F.3. Prompt for Identifying the Start POI
Prompt for Start POI Identification Please act as a top-tier AI travel planning assistant. Your job is to return the index of the best starting point for a day trip itinerary based on user needs and provided candidate POIs. If the candidate POIs are empty, please default to returning [”0”]. ### Task Overview Your task is to return the index of the best starting point for an itinerary based on the given candidate POIs and the user request. Directly return the starting points index based on user needs and candidate POIs, without writing any code. ### Candidate POIs {candidate_strings} ### User Request {user_reqs} ### Guidelines 1. Ensure the selected POI meets the user request. 2. The starting point should be close to its neighboring points. 3. Prioritize POIs like museums or art galleries, which usually require more exploration time. 4. Avoid starting from bars or clubs. ### Example Inputs and Outputs - **Example 1**: - **Candidate POIs**: [”Museum”, Park”, Bar”] - **Output**: [”0”] - **Example 2**: - **Candidate POIs**: [”Shopping Center”, Art Gallery”, Historical Building”] - **Output**: [”1”] ### Output Specification 1. Return a list of length 1, containing an integer that represents the index of the best starting point. 2. Return as a JSON List, with only one element inside. 3. The output should not contain any other information, ensuring it can be parsed by json.loads. 4. Your response should be a length-1 JSON list, where the index comes from {return_candidates}. - Example: [”0”] Now, based on the **Candidate POIs** and **User Needs**, return the index of the best starting point for the day trip itinerary according to the **Output Specification**. Note, ensure your reply is **a list composed of a single number** from {return_candidates}, following the requirements in the **Output Specification** to return **a length-1 JSON list**, and do not return any other content.
F.4. Prompt for Generating the Itinerary
Prompt for final itinerary generation You are a highly creative and knowledgeable tour guide, specifically to design a perfect day trip itinerary. Please consider carefully and use the provided Candidate POIs list to craft a one-day itinerary in the form of an engaging and realistic travel story. ## Itinerary Information Next, please follow the guidelines I provide to design a memorable day trip itinerary for tourists. Design a day trip itinerary for tourists: - **Order of candidate POIs**: {context_string} - **Must-see POIs**: {keyword_reqs} - **Keyword Requirements**: {keyword_reqs} - **Users Original Request**: {userReqList} - **Start POI**: {start_poi} - **End POI**: {start_poi} ## Constraints - **Itinerary time**: Less than {hours} hours - **POI selection**: Must follow the given sequence order ## Output Format: {{ itinerary”: List of POIs, separated by ’->’” Overall reason”: Overall recommendation reason for the designed day trip itinerary”, pois”: {{ n”: Description and recommendation reason for each POI”, }} }} Note: - n is the sequence number, which should be an integer. Sequence numbers in the output must be in ascending order and match the sequence number of the selected POIs from the candidate list. - itinerary lists all the POIs names visited, separated by ’->’, such as poi1->poi2->…”, note that it includes names only, without sequence numbers, and the order is consistent with the order of POIs in pois”. ## Pre-action Considerations 1. Work on problems step-by-step. 2. Do not omit or simplify anything. 3. Ensure the tourists feel that the itinerary is tailor-made for them. 4. **ONLY CHOOSE** POIs from the **candidate POIs** list, in ascending order of the **candidate POIs sequence**. 5. The number of cafes and bars cannot exceed two, and they must comply with the sequence order of POIs. **Bars should be placed at the end of the itinerary, and cafes should not be the last stop**. 6. Ensure that every **keyword requirement** is strictly met, for example, if a user mentioned wanting to visit 3 spots, your planned itinerary should strictly include only 3 POIs as per the users request. ## Itinerary Creation Steps 1. Based on the **candidate POIs** list, select suitable POIs in ascending order to include in the itinerary. Ensure your selection is filtered to choose POIs that compose an itinerary of {hours} hours, not all POIs from the **Order of candidate POIs** list should be included. 2. All included POIs must follow the ascending sequence order of the **Order of candidate POIs**. 3. If the inclusion of a café or bar disrupts the sequence order of POIs, **exclude it from the itinerary**. 4. Ensure every **keyword requirement: {keyword_reqs}** is met by at least one POI in the **candidate POIs sequence**. 5. **Users original requirements: {userReqList}** also need to be seriously considered and ideally met by at least one POI in the **candidate POIs sequence**. 6. Finally, generate a JSON file containing all selected POIs. Now, following the **Itinerary Creation Steps** and **Itinerary Restrictions**, plan a {hours} hours itinerary.
F.5. Prompt for Baseline
We provide the prompt for the baseline GPT-4 CoT below. For baselines without CoT, we remove the “Think step by step” part and “thoughts” in the output format.
Prompt for baseline GPT-4 CoT You are a travel planning assistant. Please help me plan a city tour itinerary in {city} based on requirements. The output should include specific Points of Interests (POI), and the itinerary should contain no less than 6 POIs: ## User Request {user_request} Think step by step: You need to understand and analyze the users request, including must-see POIs, positive requests, negative requests, etc., and then provide your recommended POIs and reasons for recommendation. ## Output Format: {{ thoughts”: Your understanding and analysis of user requirements”, itinerary”: List of POIs, separated by ’->’”, overall_reason”: The overall reason for recommending this one-day tour itinerary”, pois”: {{ n”: Description and reason for recommending the POI”, }} }} n starts from 1 and increments. Please strictly follow the output format to return JSON.
F.6. Prompt for GPT Evaluation
Prompt for GPT evaluation is provided below.
Prompt for GPT Evaluation You are a professional travel assistant. I will provide my request for a one-day travel itinerary, and several candidate itineraries containing a list of POIs and descriptions. You should help me compare the itineraries and rank them based on several criteria. ## Criteria 1. POI Quality: how interesting and diverse the POIs are 2. Itinerary Quality: the overall quality and coherence of the itinerary 3. Matchness: the matchness between the itinerary and the user query ## Input Format Each input candidate itinerary is a dictionary in the following format: {{ itinerary”: a list of POIs, separated by ’->’” overall_reason”: The overall recommendation reason for the designed one-day travel itinerary”, pois”: {{ n”: description of the POI”, }} }} ## Request {user_request} ## Candidate Itineraries {itineraries} ## Output Format Output a json object (dictionary) with four keys: poi_quality”, itinerary_quality”, matchness”, language_quality”. Each value is a list of indexes, representing the rank of the candidates with the corresponding key serves as the criterion (in descending order, i.e. from high to low). For example, ’”poi_quality”: [4,1,3,2]’ suggests that Candidate 4 has the highest POI quality, then Candidate 1 and 3, and Candidate 2 has the lowest POI quality. Ensure that your output can be parsed with Python json.loads. Do not output anything else.